OpenStudy (anonymous):

which of the following best describes the solution to the equation? 10x+20=-30+10x a. no real solutions b. exactly one real solution, x=-1and 3/4 c. infinite many real solutions d. exactly one solution,x=-1 and 1/2

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

did u solve the equation

OpenStudy (anonymous):

i did not

OpenStudy (anonymous):

well then lets do it together

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

10x+20=-30+10x +30 +30 10x+50=10x -10x -10x

OpenStudy (anonymous):

so automatically you know its not D or B

OpenStudy (anonymous):

so which one do u think it is

OpenStudy (anonymous):

no solution

OpenStudy (anonymous):

yup :)

OpenStudy (anonymous):

yea thnx